QUESTION 3 A company wants to create a market value-weighted index from a group of stocks with a total market capitalization of $600,000. If the desired initial index value is 1,500, then what is the divisor? 400 QUESTION 4 A constituent stock of a market value-weighted index splits 2-for-1. Which one of the following best describes how the stock index changes? O The total market capitalization increases, divisor decreases, but the index value does not change following the stock split. O The total market capitalization and the divisor do not change, but the index value increases following the stock split. The total market capitalization, divisor, and the index value do not change following the stock split. The total market capitalization decreases, divisor increases, but the index value does not change following the stock split. The total market capitalization and the divisor do not change, but the index value decreases following the stock split.
Answer #1

Share

Outstanding shares

(Base year)

Base price Current price

A

800 50 80
B 500 40 60
C 300 60 100

Assumed base index is 1000

Share

Outstanding shares

(Base year)

Base price Base Value Current price Current Value

A

800 50 40000 80 64000
B 500 40 20000 60 30000
C 300 60 18000 100 30000
Total Value 78000 124000
Market value-weighted index (124000/78000)*100 158.97

Divisor = 124000/1000 = 124

Suppose, stock B has spit 2 for 1. That means now total B stocks are 1000.

Share

Outstanding shares

(Base year)

Base price Base Value Current price Current Value

A

800 50 40000 80 64000
B 500 40 20000 60 60000(WN)
C 300 60 18000 100 30000
Total Value 78000 154000
Market value-weighted index (154000/78000)*100 197.44

Working Note: 60*1000(new shares) = 60000

Divisor = 154000/1000 = 154

Option 3 is the answer. That is, the

total market capitalization will increase because the number of share/stocks has doubled in amount.

Divisor will increase because the market capitalization has increased.

The index will also increase.